### 01 Jan 2024 v3

# Some Fibonacci-Related Sequences

Abstract
We discuss an interesting sequence defined recursively; namely, sequence A105774 from the On-Line Encyclopedia of Integer Sequences, and study some of its properties. Our main tools are Fibonacci representation, finite automata, and the Walnut theorem-prover. We also prove two new results about synchronized sequences.
1 Introduction
Define the Fibonacci numbers $(F_{n})_{n\geq 0}$ as usual by the initial values $F_{0}=0$, $F_{1}=1$, and $F_{n}=F_{n-1}+F_{n-2}$ for $n\geq 2$. In 2005, the first author defined an interesting sequence, A105774, $(a(n))_{n\geq 0}$ as follows:
$a(n)=\begin{cases}n,&\text{if $n\leq 1$};\\ F_{j+1}-a(n-F_{j}),&\text{if $F_{j}<n\leq F_{j+1}$ for $j\geq 2$}.\end{cases}$ (1)
The first few values of $(a(n))$ are given in Table 1. Although strictly speaking the sequence was originally defined only for $n\geq 1$, it makes sense to set $a(0)=0$, which we have done.
$n$ 0 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19 20
$a(n)$ 0 1 1 2 4 4 7 7 6 12 12 11 9 9 20 20 19 17 17 14 14
The sequence has an intricate fractal structure, which is depicted in Figure 1.
Figure 1: Graph of $a(n)$.
A number of properties of this sequence were stated, without proof, in the index entry for A105774 in the On-Line Encyclopedia of Integer Sequences (OEIS) [11]. In this paper we prove these properties, and many new ones, with the aid of finite automata and the Walnut theorem-prover [8, 10].
We will need the Lucas numbers, $L_{n}$, defined by $L_{0}=2$, $L_{1}=1$, and $L_{n}=L_{n-1}+L_{n-2}$ for $n\geq 2$. Recall the Binet forms:
$F_{n}$ $=(\varphi^{n}-(-1/\varphi)^{n})/\sqrt{5}$ (2)
$L_{n}$ $=\varphi^{n}+(-1/\varphi)^{n}.$ (3)
In addition to studying A105774, we also prove two new results about synchronized sequences, in Theorems 10 and 18.
2 Fibonacci representations
In this paper we represent the natural numbers =${\mathbb{N}}=\{0,1,2,\ldots\}$ in the Fibonacci (aka Zeckendorf) numeration system [7, 12]. In this system we write $n$ as the sum of distinct Fibonacci numbers $F_{i}$ for $i\geq 2$, subject to the condition that no two consecutive Fibonacci numbers appear in the sum. For example, $143=F_{9}+F_{6}+F_{2}=34+8+$ 1. Usually we write this representation as a bit string $(n)_{F}=e_{1}e_{2}\cdots e_{t}$, with $e_{i}\in\{0,1\}$, $e_{1}\not=0$, and $e_{i}e_{i+1}\not=1$, such that $n=\sum_{1\leq i\leq t}e_{i}F_{t-i+2}$. For example, $(43)_{F}=10010001$. The inverse function, mapping an arbitrary bit string (with no conditions) $x=e_{1}e_{2}\cdots e_{t}$ to the number it represents is defined to be $[x]_{F}=\sum_{1\leq i\leq t}e_{i}F_{t-i+2}$.
3 Finite automata and regular expressions
A finite automaton is a simple model of a computer. In its simplest incarnation, a deterministic finite automaton (DFA), the automaton can be in any one of a finite number of different states. It takes as input a finite string over a finite alphabet, and processes its input symbol-by-symbol, choosing the next state through a lookup table (the transition function) based only on the current state and the next input symbol. Some of the states are designated as accepting states; an input string is accepted if the automaton is in an accepting state after all its symbols are processed. In this paper we represent integers in Fibonacci representation, so we can think of the DFA as processing integers instead of strings. Thus, a DFA is a good model to represent a set of natural numbers; namely, the set of all integers whose representations are accepted. In this sense, a DFA decides membership for a set (or sequence). DFA’s are typically drawn as transition diagrams, where the arrows represent transitions between states, double circles denote accepting states, and single circles denote nonaccepting states.
As an example, consider the automaton in Figure 2. It takes the Fibonacci representation of a natural number $n$ as input, and accepts if and only if $n$ is even.
Figure 2: Fibonacci automaton for the set of even numbers.
A slightly more complicated model is the deterministic finite automaton with output (DFAO). In this model, an output is associated with every state, and the output on an input string $x$ is the output associated with the last state reached after processing $x$. This model is adequate for representing sequences over a finite alphabet; since there are only finitely many states, only finitely many different outputs are possible. If a sequence $(a(n))$ is computed by a DFAO in the following way—the input is the representation of $n$ and the output is $a(n)$—then we say $(a(n))$ is automatic. Again, a DFAO is often displayed as a transition diagram, with the notation $q/i$ in a state meaning that the state is numbered $q$ and has output $i$.
For example, the transition diagram in Figure 3 illustrates a DFAO computing the sequence $(n\bmod 3)_{n\geq 0}$, where the input is in Fibonacci representation.
Figure 3: Fibonacci DFAO computing $n\bmod 3$.
We note in passing that, due to a nice result of Charlier et al. [4], the minimal DFAO computing $(n\bmod k)_{n\geq 0}$ for $n$ in Fibonacci representation is known to have $2k^{2}$ states.
Finally, there is the notion of synchronized automaton that computes a function $f:{\mathbb{N}}\rightarrow{\mathbb{N}}$. In this case, we use a DFA, but with two integer inputs $n$ and $x$ read in parallel, where the shorter input is padded with leading zeros, if necessary, and the DFA accepts if and only if $x=f(n)$. This model is suited to representing sequences of natural numbers. It has the important advantage that if a sequence is represented in this way, then there is a decision procedure to answer first-order queries about the values of the sequence, in the logical structure $\langle{\mathbb{N}},+,n\rightarrow f(n)\rangle$. See [9] for more details.
As an example, the synchronized automaton in Figure 4 computes the function $n\rightarrow\lfloor n/2\rfloor$ in Fibonacci representation.
Figure 4: Synchronized Fibonacci automaton computing $\lfloor n/2\rfloor$.
It is quite useful to adopt the convention that in all three models, the result is not sensitive to the presence or absence of leading zeros in the representation of numbers.
The most powerful of all three representations for a sequence is the synchronized automaton, since one can obtain the other two from it easily. The downside is that it is quite possible that there exists a DFA to decide membership in a sequence $\bf s$, but there is no corresponding synchronized automaton computing the $n$’th term of $\bf s$.
In this paper, we sometimes use the notation called a regular expression. For our purposes, the main thing to know is that $x^{*}$, where $x$ is a single string or set of strings, denotes the set of all strings consisting of zero or more concatenations of elements chosen from the set $x$ specifies. For more about automata and regular expressions, see [6].
4 Walnut
Walnut is a theorem-prover for automatic sequences. It can answer first-order queries about automatic and synchronized sequences defined over the natural numbers ${\mathbb{N}}$. The syntax is basically first-order logic, with the following meanings of symbols:
  • & is logical AND, | is logical OR, ~ is logical NOT, => is implication, and <=> means IFF.
  • != means $\neq$ and x/y denotes integer division, that is, $\lfloor x/y\rfloor$.
  • A is Walnut’s representation for the universal quantifier $\forall$ and E is the representation for $\exists$.
  • reg defines a regular expression.
  • def defines an automaton for later use.
  • eval evaluates a first-order logic expression with no free variables, and returns either TRUE or FALSE.
  • ?msd_fib is a bit of jargon telling Walnut to represent all numbers in the Fibonacci numeration system.
When Walnut returns TRUE (resp., FALSE) to an eval query, the result is a rigorous proof that the particular logical statement holds (resp., does not hold).
As examples, here is the Walnut code for generating the automata in Figures 2 and 4:
def even "?msd_fib Ek n=2*k":
def nover2 "?msd_fib z=n/2":
We will need some Walnut code for the functions $n\rightarrow\lfloor\varphi n\rfloor$ and $n\rightarrow\lfloor\varphi^{2}n\rfloor$, where $\varphi=(1+\sqrt{5})/2$, the golden ratio. The synchronized automata for these can be found in [10], under the names phin and phi2n, respectively.
5 Guessing and checking
One of the principal tools we use is the “guess-and-check” method. Here we guess a finite automaton computing a sequence based on empirical data, and then use Walnut to verify that it is correct, usually by checking that the defining equation holds. The guessing procedure is described in [10, §5.7].
We can carry out this guessing procedure for the sequence A105774, which we call $(a(n))$. It produces a synchronized automaton that computes the sequence in the following sense: the input is a pair $(n,x)$ in Fibonacci representation, and the automaton accepts if and only if $x=a(n)$. The automaton a105774 that we guess is displayed in Figure 5.
Figure 5: Synchronized Fibonacci automaton for $a(n)$.
The next step, which is crucial, is to verify that our guessed automaton is actually correct. First we must verify that the automaton really does compute a function. This means that for each $n$ there is exactly one $x$ such that the pair $(n,x)$ is accepted. We can verify this as follows:
eval check_at_least_one "?msd_fib An Ex \$a105774(n,x)":
eval check_at_most_one "?msd_fib ~En,x1,x2 x1!=x2 & \$a105774(n,x1) &
   \$a105774(n,x2)":
and Walnut returns TRUE for both.
Next we must verify that our automaton obeys the defining recurrence (1).
reg adjfib msd_fib msd_fib "[0,0]*[0,1][1,0][0,0]*":
[…]
def trapfib "?msd_fib \$adjfib(x,y) & x<k & y>=k":
# accepts (k,x,y) if x is the largest Fibonacci number
# less than k and y is the next largest Fib number
eval test105774 "?msd_fib Ak,x,y,z,t (\$trapfib(k,x,y) & \$a105774(k,z)
   & \$a105774(k-x,t)) => y=z+t":
and Walnut returns TRUE. At this point we know that our guessed automaton is correct.
Let’s start by proving a basic property of A105774.
Proposition 1.
No natural number appears three or more times in A105774.
Proof.
We use the following Walnut code.
eval test012 "?msd_fib ~Ex,y,z,n x<y & y<z & \$a105774(x,n) &
   \$a105774(y,n) & \$a105774(z,n)":
and Walnut returns TRUE. ∎
Now let’s create a DFAO computing $c(n)$, the number of times each natural number appears in A105774. We can do this with the following Walnut commands:
def s0 "?msd_fib ~Ex \$a105774(x,n)":
def s2 "?msd_fib Ex,y x<y & \$a105774(x,n) & \$a105774(y,n)":
def s1 "?msd_fib ~(\$s0(n)|\$s2(n))":
combine C s1=1 s2=2 s0=0:
This creates the DFAO in Figure 6.
Figure 6: Fibonacci DFAO for $c(n)$.
The first few values of the sequence $c(n)$ are given in Table 5.
$n$ 0 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19 20
$c(n)$ 1 2 1 0 2 0 1 2 0 2 0 1 2 0 2 1 0 2 0 1 2
It is sequence A368199 in the OEIS.
Proposition 2.
If $n$ appears twice in A105774, the two occurrences are consecutive.
Proof.
We use the following Walnut code:
eval twice_consec "?msd_fib An,x,y (x<y & \$a105774(x,n) & \$a105774(y,n))
   => y=x+1":
and Walnut returns TRUE. ∎
The positions $p_{0}(n)$ of the $n$’th 00 (respectively, the $n$’th $1$ and the $n$’th $2$) in A368199 are also Fibonacci-synchronized. Note that we index starting at $n=0$. Once again, we can prove this by guessing and checking. We omit the checks that our guessed automata p0,p1,p2 compute synchronized functions.
eval chek1a "?msd_fib An C[n]=@1 <=> Ek \$p1(k,n)":
eval chek2a "?msd_fib An C[n]=@2 <=> Ek \$p2(k,n)":
eval chek0b "?msd_fib Aj,m,n (\$p0(j,m) & \$p0(j+1,n)) => m<n":
eval chek1b "?msd_fib Aj,m,n (\$p1(j,m) & \$p1(j+1,n)) => m<n":
eval chek2b "?msd_fib Aj,m,n (\$p2(j,m) & \$p2(j+1,n)) => m<n":
The first few values of $p_{0}$, $p_{1}$, and $p_{2}$ are given in Table 5.
$n$ 0 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18
$p_{0}(n)$ 3 5 8 10 13 16 18 21 24 26 29 31 34 37 39 42 45 47 50
$p_{1}(n)$ 0 2 6 11 15 19 23 28 32 36 40 44 49 53 57 61 66 70 74
$p_{2}(n)$ 1 4 7 9 12 14 17 20 22 25 27 30 33 35 38 41 43 46 48
Each of these sequences is already in the OEIS. We now prove the characterizations.
Proposition 3.
The numbers $p_{2}(n)$ are precisely those in A007064, that is, the natural numbers not of the form $\lfloor\varphi k+{1\over 2}\rfloor$ for $k\geq 0$.
Proof.
The sequence A007067 is defined to be $\lfloor n\varphi+1/2\rfloor$. Now
$\lfloor n\varphi+1/2\rfloor=\left\lfloor{{2n\varphi+1}\over 2}\right\rfloor\\ =\left\lfloor{{\lfloor 2n\varphi+1\rfloor}\over 2}\right\rfloor\\ =\left\lfloor{{\lfloor 2n\varphi\rfloor+1}\over 2}\right\rfloor,$
where we have used the fact that $\lfloor{{\lfloor x\rfloor}\over n}\rfloor=\lfloor x/n\rfloor$ for real numbers $x$ and integers $n\geq 1$. Hence A007067 can be computed by the following Walnut code:
def a007067 "?msd_fib Ex \$phin(2*n,x) & z=(x+1)/2":
The sequence A007064 consists of those integers not in A007067. It is not hard to see that A007064 is given (after a change of index) by $n\rightarrow\lfloor n\varphi+\varphi/2\rfloor+n+1$ for $n\geq 0$. Indeed, we can verify this as follows:
def a007067 "?msd_fib Ex \$phin(2*n,x) & z=(x+1)/2":
def a007064 "?msd_fib Ex \$phin(2*n+1,x) & z=n+1+x/2":
eval check_two "?msd_fib Ax (Em \$a007067(m,x)) <=> (~En \$a007064(n,x))":
Finally we can prove the result about $p_{2}$ as follows:
eval checkp2 "?msd_fib An (Ek \$p2(k,n)) <=> (Em \$a007064(m,n))":
and Walnut returns TRUE. ∎
Proposition 4.
The numbers $p_{1}(n)$, $n\geq 1$, are precisely those integers in A035487.
Proof.
We use the following Walnut command:
def a035487 "?msd_fib Ex \$a007064(x) & \$a007067(x,n)":
eval checkp1 "?msd_fib An (n>0) => (\$a035487(n) <=> (Ek \$p1(k,n)))":
and Walnut returns TRUE. ∎
Finally, we prove the result for $p_{0}$:
Proposition 5.
The natural numbers $p_{0}(n)$, i.e., the numbers that do not appear in A105774 are those given by the sequence A004937, that is, $\{\lfloor\varphi^{2}n+{1\over 2}\rfloor\,:\,n\geq 1\}$.
Proof.
We have, using the same idea as in the proof of Proposition 3, that
$\lfloor\varphi^{2}n+1/2\rfloor=\left\lfloor{{{\lfloor\varphi^{2}2n\rfloor}+1}\over 2}\right\rfloor.$
We can now use Walnut to define a synchronized automaton for A004937:
def a004937 "?msd_fib Ex \$phi2n(2*n,x) & z=(x+1)/2":
The automaton for A004937 is given in Figure 7.
Figure 7: Synchronized Fibonacci DFAO for A004937.
We can now complete the proof as follows:
eval chk0 "?msd_fib An (n>0) => ((Ek \$a004937(k,n)) <=> (Ej \$p0(j,n)))":
and Walnut returns TRUE. ∎
6 Inequalities
Proposition 6.
For all $n\geq 0$ we have $\lfloor{{\varphi+2}\over 5}n\rfloor\leq a(n)\leq\lfloor\varphi n\rfloor$.
Proof.
For the first inequality we write
eval lowerbound "?msd_fib An,x,y (\$a105774(n,x) & \$phin(n,y)) => x>=(y+2*n)/5":
For the second inequality we use the following Walnut code:
eval upperbound "?msd_fib An,x,y (\$a105774(n,x) & \$phin(n,y)) => x<=y":
and Walnut returns TRUE. ∎
The bounds in Proposition 6 are tight, as the following theorem shows.
Proposition 7.
We have $\liminf_{n\rightarrow\infty}a(n)/n={{\varphi+2}\over 5}$ and $\limsup_{n\rightarrow\infty}a(n)/n=\varphi$.
Proof.
For the first claim, using the well-known Binet formulas for the Fibonacci and Lucas numbers, it suffices to show that $a(L_{k}+1)=F_{k+1}+1$ for all $k\geq 3$.
reg lucfib msd_fib msd_fib "[0,0]*[1,1][0,0][1,0][0,0]*":
[…]
eval chklow "?msd_fib Ax,y \$lucfib(x,y) => \$a105774(x+1,y+1)":
For the second result it suffices to show that $a(F_{k}+1)=F_{k+1}-1$ for all $k\geq 2$. This follows directly from the defining recurrence for $a(n)$. Or one can use Walnut:
eval chkup "?msd_fib Ax,y,m (\$adjfib(x,y) & \$a105774(x+1,m)) => m+1=y":
The graph in Figure 1 suggests studying what the positions of what might be called “suffix minima”: those $n$ for which $a(m)>a(n)$ for all $m>n$.
Theorem 8.
The suffix minima of $(a(n))$ for $n>0$ occur precisely when $(n)_{F}\in 10(100^{*}10)^{*}0^{*}$.
Proof.
We use the following Walnut code:
def suffmin "?msd_fib Am,x,y (m>n & \$a105774(m,x) & \$a105774(n,y)) => x>y":
The resulting automaton is depicted in Figure 8, from which the regular expression is easily read off.
Figure 8: Fibonacci representation of positions of suffix minima.
7 Consecutive identical or different terms
Proposition 9.
We have
  • (a)
    $a(n)=a(n+1)$ if and only if $n\in\{\lfloor\varphi^{2}k\rfloor-1\,:\,k\geq 1\}$;
  • (b)
    $a(n)\not=a(n+1)$ if and only if $n\in\{\lfloor\varphi k+1/\varphi\rfloor\,:\,k\geq 0\}$.
  • (c)
    $a(n)\not\in\{a(n-1),a(n+1)\}$ if and only if $n\in\{\lfloor\varphi\lfloor\varphi^{2}k\rfloor\rfloor\,:\,k\geq 1\}$.
Proof.
  • (a)
    We use the following Walnut command:
    eval twoconsec "?msd_fib An (Ex \$a105774(n,x) & \$a105774(n+1,x)) <=>
       (Ek,y (k>0) & \$phi2n(k,y) & y=n+1)":
  • (b)
    We use the following Walnut command:
    eval differ "?msd_fib An (Ex,y \$a105774(n,x) & \$a105774(n+1,y) &
       x!=y) <=> (Ek,y \$phin(k+1,y) & y=n+1)":
  • (c)
    We use the following Walnut commands:
    def a003623 "?msd_fib Ex \$phi2n(n,x) & \$phin(x,z)":
    eval isolated "?msd_fib An (n>0) => ((Ex,y,z \$a105774(n-1,x) &
       \$a105774(n,y) & \$a105774(n+1,z) & x!=y & y!=z) <=>
       (Ek \$a003623(k,n)))":
The automaton computing A003623 is given in Figure 9.
Figure 9: Synchronized Fibonacci DFAO for A003623.
8 Sorting the terms
We can consider sorting the terms of A105774 in ascending order. The resulting sequence is A368200. We can now guess a Fibonacci automaton for this sequence; it is displayed in Figure 10.
Figure 10: Synchronized Fibonacci DFAO for A368200.
It now remains to verify that this is indeed A105774 in ascending order. This is a consequence of a more general (and new) result, as follows.
Theorem 10.
Suppose $(a(n))_{n\geq 0}$ and $(b(n))_{n\geq 0}$ are synchronized sequences. Then the following property is decidable: $(a(n))_{n\geq 0}$ is a permutation of $(b(n))_{n\geq 0}$
Proof.
We just sketch the proof, which depends on linear representations. For more about these, see [10, Chap. 9] and [2]. Given automata for $(a(n))_{n\geq 0}$ and $(b(n))_{n\geq 0}$, we can find linear representations counting the number of occurrences of each natural number $n$ in each sequence. Then we can form the linear representation for the difference of these two. Then we use the fact that whether a linear representation is identically 00 is decidable. ∎
When we perform this calculation for the sequences A105774 and A368200, we find that the linear representation for the difference is indeed zero. It now simply remains to check that the numbers computed by the automaton are in ascending order:
eval ascending "?msd_fib An,x,y (\$a368200(n,x) & \$a368200(n+1,y)) => y >= x":
Proposition 11.
[…]
Proof.
We use the following Walnut code.
def diff "?msd_fib Ex,y \$a368200(n,x) & \$a368200(n+1,y) & y=x+z":
def checkdiff "?msd_fib An,z \$diff(n,z) => (z=0|z=1|z=2)":
It now follows that the first difference sequence $(d(n))_{n\geq 0}$ defined by $d(n)=b(n+1)-b(n)$ is Fibonacci-automatic. In fact, it is $c(n)$ in disguise, as the next theorem shows.
Theorem 12.
For $n\geq 0$ we have $d(n)=2-c(n)$.
Proof.
We use the following Walnut commands:
eval cd0 "?msd_fib An \$diff(n,0) <=> C[n]=@2":
eval cd1 "?msd_fib An \$diff(n,1) <=> C[n]=@1":
eval cd1 "?msd_fib An \$diff(n,2) <=> C[n]=@0":
and Walnut returns TRUE each time. ∎
9 Special values
In this section we compute the values of $a(n)$ when $n$ is a Fibonacci or Lucas number.
Theorem 13.
Let $s(n)=a(F_{n})$.
  • (a)
    The Fibonacci representation of $s(n)$ for $n\geq 3$ is $(1000)^{\lfloor(n-3)/4\rfloor}x_{n}$ where
    $x_{n}=\begin{cases}10,&\text{if $n\equiv 0\ \mbox{\rm(mod}\ 4\mbox{\rm)}$;}\\ 101,&\text{if $n\equiv 1\ \mbox{\rm(mod}\ 4\mbox{\rm)}$;}\\ 1001,&\text{if $n\equiv 2\ \mbox{\rm(mod}\ 4\mbox{\rm)}$;}\\ 1,&\text{if $n\equiv 3\ \mbox{\rm(mod}\ 4\mbox{\rm)}$.}\end{cases}$
  • (b)
    $(s(n-1))$ is the OEIS sequence A006498, satisfying the recurrence $s(n)=s(n-1)+s(n-3)+s(n-4)$ for $n\geq 4$. A closed form is as follows:
    $s(n)=L_{n}/10+F_{n}/2+\begin{cases}-{1\over 5}(-1)^{n/2},&\text{$n$ even;}\\ {2\over 5}(-1)^{(n-1)/2},&\text{$n$ odd.}\end{cases}$
Proof.
For part (a) we create a synchronized automaton that accepts the Fibonacci representation of $F_{n}$, which is $10^{n-2}$, in parallel with $s(n)$. It is displayed in Figure 11, from which the result immediately follows by inspection.
reg isfib msd_fib "0*10*":
def special "?msd_fib \$isfib(x) & \$a105774(x,y)":
Figure 11: Synchronized Fibonacci DFAO for $s(n)$.
For part (b) we first create a DFA that accepts, in parallel, the Fibonacci representations of $F_{n+4},F_{n+3},F_{n+1},F_{n}$. Then we check the recurrence.
reg four msd_fib msd_fib msd_fib msd_fib
   "[0,0,0,0]*[1,0,0,0][0,1,0,0][0,0,0,0][0,0,1,0][0,0,0,1][0,0,0,0]*":
eval partb "?msd_fib Aa,b,c,d,x,y,z,w (\$four(a,b,c,d) & \$a105774(a,x) &
   \$a105774(b,y) & \$a105774(c,z) & \$a105774(d,w)) => x=y+z+w":
and Walnut returns TRUE. ∎
Proposition 14.
  • (a)
    Suppose $k\geq 3$. Over the range $F_{k}\leq n<F_{k+1}$, the sequence $a(n)$ achieves its minimum uniquely at $n=F_{k}$.
  • (b)
    Suppose $k\geq 5$. Over the range $F_{k}\leq n<F_{k+1}$, the sequence $a(n)$ achieves its maximum only at $n=F_{k}+1$ and $n=F_{k}+2$.
Proof.
We use the following Walnut code:
eval minval "?msd_fib Ax,y,z,t,u (\$adjfib(x,y) & \$a105774(x,z) &
   t>x & t<y & \$a105774(t,u)) => u>z":
eval maxval "?msd_fib Ax,y,z,t,u,w (x>=5 & \$adjfib(x,y) & \$a105774(x+1,z) &
\$a105774(x+2,w) & t>=x & t<y & \$a105774(t,u)) => (z=w & u<=z)":
and Walnut returns TRUE for both. ∎
Corollary 15.
Suppose $F_{k}\leq n<F_{k+1}$. Then $s(k)\leq a(n)\leq F_{k+1}-1$.
Theorem 16.
Define $t(n)=a(L_{n})$ for $n\geq 0$. Then $t(n)=t(n-1)+t(n-3)+t(n-4)$ for $n\geq 5$. A closed form for $n\geq 2$ is given by
$t(n)=3L_{n}/10+3F_{n}/2+\begin{cases}{2\over 5}(-1)^{n/2},&\text{$n$ even;}\\ {1\over 5}(-1)^{(n-1)/2},&\text{$n$ odd.}\end{cases}$
Proof.
Same ideas as in Theorem 13. We omit the details. ∎
10 Parity
Proposition 17.
[…]
Proof.
We use the following Walnut code:
def even "?msd_fib Ek n=2*k":
eval checkparity "?msd_fib An,x,y (\$a105774(n,x) & \$phin(n,y)) =>
   (\$even(x) <=> \$even(y))":
and Walnut returns TRUE. ∎
11 Distinct values
Given a sequence taking values in ${\mathbb{N}}$ we can consider the same sequence obtained by scanning the terms from left-to-right, and deleting any values that previous occurred in the sequence. We call this the distinctness transform of the original sequence. For example, the distinctness transform of A105774 starts $0,1,2,4,7,6,12,11,\ldots$. We have the following new theorem.
Theorem 18.
The following problem is decidable: given two synchronized automata computing $s(n)$ and $s^{\prime}(n)$, whether $(s^{\prime}(n))$ is the distinctness transform of $(s(n))$.
Proof.
It suffices to create a first-order logical formula asserting that $(s^{\prime}(n))$ is the distinctness transform of $(s(n))$.
This formula asserts
  • (a)
    An integer occurs in the range of $s(n)$ iff it occurs in the range of $s^{\prime}(n)$;
  • (b)
    All the values $s^{\prime}(n)$ for $n\geq 0$ are distinct;
  • (c)
    For all $x$ and $y$, if $x$’s first occurrence in $(s(n))$ precedes that of $y$’s first occurrence, then $x$’s unique occurrence in $(s^{\prime}(n))$ also precedes that of $y$.
Let us carry this out for A105774. Let $(a^{\prime}(n))$ denote the distinctness transform of $(a(n))$. The first few values are given in Table 11.
$n$ 0 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19
$a^{\prime}(n)$ 0 1 2 4 7 6 12 11 9 20 19 17 14 15 33 32 30 27 28 22
We can guess the automaton for this sequence and then verify it.
Figure 12: Synchronized Fibonacci automaton for $a^{\prime}(n)$.
First let us check that our guessed automaton really does compute a function:
def checkap1 "?msd_fib An Ex \$aprime(n,x)":
def checkap2 "?msd_fib ~En,x1,x2 x1!=x2 & \$aprime(n,x1) & \$aprime(n,x2)":
and Walnut returns TRUE for both.
Next, let’s check the three conditions guaranteeing that $a^{\prime}$ is the distinctness transform of $a$:
eval check_distinct1 "?msd_fib Ax (Em \$a105774(m,x)) <=> (En \$aprime(n,x))":
eval check_distinct2 "?msd_fib ~En1,n2,x n1!=n2 & \$aprime(n1,x) &
   \$aprime(n2,x)":
def first_occ "?msd_fib \$a105774(y,n) & Ax (x<y) => ~\$a105774(x,n)":
# provided n occurs in A105774, gives the least position y
# where it occurs
eval check_distinct3 "?msd_fib Ax,y,i,j (\$first_occ(x,i) & \$first_occ(y,j)
   & i<j) => Em,n \$aprime(m,x) & \$aprime(n,y) & m<n":
and Walnut returns TRUE for all of them.
We now examine the run-length encoding of A105774. The run-length encoding of an arbitrary sequence $(s(n))_{n\geq 0}$ counts the size of maximal blocks of consecutive identical values [5]. For A105774 it is $1,2,1,2,2,1,2,1,2,\ldots$.
Once we have the distinct terms of A105774, we can easily determine the run-length encoding of the sequence.
Theorem 19.
The run-length encoding of A105774 is A001468, that is,
$1(2-{\bf f}[0])(2-{\bf f}[1])(2-{\bf f}[2])\cdots$
where ${\bf f}=01001\cdots$ is the infinite Fibonacci word [1].
Proof.
The claim can be verified by the following Walnut code:
def nthrun2 "?msd_fib Ex,y \$aprime(n,x) & \$first_occ(x,y) &
   \$a105774(y+1,x)":
eval compare_fib "?msd_fib An \$nthrun2(n+1) <=> F[n]=@0":
12 Other results
Define $w(n)$ to be the least integer such that $a(w(n))\geq n$. The first few terms of this sequence are given in Table 12.
$n$ 0 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19 20
$w(n)$ 0 1 3 4 4 6 6 6 9 9 9 9 9 14 14 14 14 14 14 14 14
Proposition 20.
If $n\geq 2$ and $F_{k}\leq n<F_{k+1}$ then $w(n)=F_{k}+1$.
Proof.
We use the following Walnut commands:
def trapfib2 "?msd_fib \$adjfib(x,y) & x<=k & y>k":
def wseq "?msd_fib (Em \$a105774(x,m) & m>=n) &
   (Ai,p (i<x & \$a105774(i,p)) => p<n)":
eval propw "?msd_fib Ax,y,n,m (n>=2 & \$trapfib2(n,x,y) & \$wseq(n,m))
   => m=x+1":
and Walnut returns TRUE. ∎
We now turn to describing the fixed points of $a$.
Theorem 21.
We have $a(n)=n$ for $n>0$ if and only if $(n)_{F}\in 1(00100^{*}1)^{*}\{\epsilon,01,010,0100\}$.
Proof.
We use the Walnut command
def fixed "?msd_fib \$a105774(n,n)":
and it produces the automaton in Figure 13, from which we can directly read off the result.
Figure 13: Solutions to $a(n)=n$.
For the remainder of this section, let $a(n)$ represent (as before) the $n$’th term of A105774, let $b(n)=\lfloor n\varphi\rfloor$, and let $c(n)=\lfloor n\varphi^{2}+1/2\rfloor$. We will find some relations concerning the composition of the functions $a$, $b$, and $c$. We introduce the function $x(n):=a(b(n))-b(a(n))$.
Theorem 22.
  • (a)
    $a(b(n))\geq b(a(n))$ for all $n\geq 0$.
  • (b)
    For $n\geq 1$, the difference $x(n):=a(b(n))-b(a(n))$ is exactly the parity of the number of $1$’s in the Fibonacci representation of $n-1$, that is, sequence A095076.
  • (c)
    $b(a(b(n)))\geq a(b(a(n)))$ for all $n\geq 0$.
  • (d)
    $a(a(b(n)))-a(b(a(n)))\in\{-2,0,2\}$ for all $n\geq 0$.
  • (e)
    If $d(n):=c(a(n))-a(b(n))-a(n)$, then $d(n)\in\{-1,0,1\}$ for all $n\geq 0$.
  • (f)
    For $n\geq 1$ we have $d(n)=0$ iff ${\bf f}[n-1]=1$.
  • (g)
    For $n\geq 1$ we have $c(a(b(n)))+a(b(n))-a(b(b(n)))-2b(a(n))=1$.
Proof.
We use the following Walnut commands:
reg even1 msd_fib "(0*10*1)*0*":
# even number of 1’s in Fibonacci representation
def ab "?msd_fib Ex \$phin(n,x) & \$a105774(x,z)":
# z = a105774(floor(phi*n))
def ba "?msd_fib Ex \$a105774(n,x) & \$phin(x,z)":
# z = floor(phi*a105774(n))
# check that ab >= ba for all n so subtraction makes sense over N:
def test "?msd_fib An,x,y (\$ab(n,x) & \$ba(n,y)) => x>=y":
# Walnut says TRUE
def xx "?msd_fib Ex,y \$ab(n,x) & \$ba(n,y) & z=x-y":
# 12 states
#test that a(b(n)) = b(a(n)) precisely when parity of 1’s in
#Fibonacci expansion of n-1 is 0:
eval test1 "?msd_fib An \$xx(n+1,0) <=> \$even1(n)":
# Walnut says TRUE
#so it follows immediately
# that a(b(n)) > b(a(n)) precisely when parity of 1’s in
#Fibonacci expansion of n-1 is 1, no need to test it
def aba "?msd_fib Ex \$ba(n,x) & \$a105774(x,z)":
# 38 states
def bab "?msd_fib Ex \$ab(n,x) & \$phin(x,z)":
# 18 states
# check the inequality bab(n) >= aba(n):
eval test3 "?msd_fib An,x,y (\$bab(n,x) & \$aba(n,y)) => x>=y":
# Walnut says true
def aab "?msd_fib Ex \$ab(n,x) & \$a105774(x,z)":
# 34 states
# check the proposition that aab(n) - aba(n) takes values in
# {-2,0,2} only:
eval test4 "?msd_fib An,x,y (\$aab(n,x) & \$aba(n,y)) => (x=y|x=y+2|y=x+2)":
# Walnut says TRUE
def a004937 "?msd_fib Ex \$phi2n(2*n,x) & z=(x+1)/2":
def ca "?msd_fib Ex \$a105774(n,x) & \$a004937(x,z)":
def dp "?msd_fib Ew,x,y \$ca(n,w) & \$ab(n,x) & \$a105774(n,y) & z+x+y=w+1":
# here dp computes d(n)+1
eval test1 "?msd_fib An,x \$dp(n,x) => (x=0|x=1|x=2)":
eval test2 "?msd_fib An (n>=1) => (F[n-1]=@1 <=> \$dp(n,1))":
def cab "?msd_fib Ex \$ab(n,x) & \$a004937(x,z)":
def abb "?msd_fib Ex \$phin(n,x) & \$ab(x,z)":
eval test3 "?msd_fib An,r,s,t,u (n>=1 & \$cab(n,r) & \$abb(n,s) & \$ab(n,t)
   & \$ba(n,u)) => r+t=s+2*u+1":
13 Carlitz-style results
[…]
We adopt the following notation: if $u=u_{1}\cdots u_{t}\in\{a,b,d,x\}^{*}$, then by $u(n)$ we mean $(u_{1}\circ\cdots\circ u_{t})(n)$. Recall that $x$ was defined in Section 12.
Theorem 23.
Let $u\in\{b,d\}^{+}$ be a nonempty word with $i$ copies of $b$ and $j$ copies of $d$. Then for all $n\geq 1$ we have
$au(n)=F_{i+2j}\cdot ab(n)+F_{i+2j-1}\cdot a(n)+C(u)\cdot(2x(n)-1),$
where $C(u)$ is defined as follows:
$C(u)=\begin{cases}0,&\text{if $u=b$};\\ 1,&\text{if $u=d$};\\ F_{i+2j-1}+C(v),&\text{if $u=vb$};\\ F_{i+2j-1}-C(v),&\text{if $u=vd$}.\end{cases}$ (4)
Proof.
First we need some lemmas:
Lemma 24.
We have
  • (a)
    $xb(n)=x(n)$ for all $n\geq 0$.
  • (b)
    $xd(n)=1-x(n)$ for all $n\geq 1$.
  • (c)
    $abb(n)=ab(n)+a(n)+2x(n)-1$ for all $n\geq 1$.
  • (d)
    $abb(n)=ad(n)$ for all $n\geq 0$.
  • (e)
    $abd(n)=2ab(n)+a(n)+2x(n)-1$ for all $n\geq 1$.
Proof.
These can be proved with Walnut as follows. Recall that phin is Walnut’s name for $b$ and phi2n is Walnut’s name for $d$.
eval checka "?msd_fib An,y,z,w (n>=0 & \$xx(n,y) & \$phin(n,z) & \$xx(z,w))
   => w=y":
eval checkb "?msd_fib An,y,z,w (n>=1 & \$xx(n,y) & \$phi2n(n,z) & \$xx(z,w))
   => w+y=1":
eval checkc "?msd_fib An,y,z,w,t (n>=1 & \$abb(n,y) & \$ab(n,z) &
   \$a105774(n,w) & \$xx(n,t)) => y+1=z+w+2*t":
def ad "?msd_fib Ew \$phi2n(n,w) & \$a105774(w,z)":
def abd "?msd_fib Ew,y \$phi2n(n,w) & \$phin(w,y) & \$a105774(y,z)":
eval checkd "?msd_fib An,z (n>=0 & \$ad(n,z)) => \$abb(n,z)":
eval checke "?msd_fib An,y,z,w,t (n>=1 & \$abd(n,y) & \$ab(n,z) &
   \$a105774(n,w) & \$xx(n,t)) => y+1=2*z+w+2*t":
and Walnut returns TRUE for all. ∎
Now we can prove the desired result by induction on $|u|$. It is clear from the above results for $|u|=1$. Now assume $|u|\geq 2$ and consider two cases: $u=vb$ and $u=vd$.
$u=vb$: In this case we have, by induction, that
$av(n)=F_{i+2j-1}\cdot ab(n)+F_{i+2j-2}\cdot a(n)+C(v)\cdot(2x(n)-1).$
Substituting $b(n)$ for $n$, we get
$avb(n)$ $=F_{i+2j-1}\cdot abb(n)+F_{i+2j-2}\cdot ab(n)+C(v)\cdot(2xb(n)-1)$
$=F_{i+2j-1}\cdot(ab(n)+a(n)+2x(n)-1)+F_{i+2j-2}ab(n)+C(v)\cdot(2xb(n)-1)$
$=F_{i+2j}\cdot ab(n)+F_{i+2j-1}a(n)+(C(v)+F_{i+2j-1})(2x(n)-1),$
as desired.
$u=vd$: By induction we have
$av(n)=F_{i+2j-2}\cdot ab(n)+F_{i+2j-3}\cdot a(n)+C(v)\cdot(2x(n)-1).$
Substituting $d(n)$ for $n$, we get
$avd(n)$ $=F_{i+2j-2}\cdot abd(n)+F_{i+2j-3}\cdot ad(n)+C(v)\cdot(2xd(n)-1)$
$=F_{i+2j-2}\cdot(2ab(n)+a(n)+2x(n)-1)+F_{i+2j-3}(ab(n)+a(n)+2x(n)-1)$
$\quad+C(v)(2(1-x(n))-1)$
$=F_{i+2j}\cdot ab(n)+F_{i+2j-1}a(n)+(F_{i+2j-1}-C(v))(2x(n)-1).$
This completes the proof. ∎
We can obtain a linear representation for $C(u)$, as follows.
Corollary 25.
Define vectors $L,R$ and a matrix-valued morphism $\mu$ as follows:
$L$ $=\left[\begin{array}[]{ccc}0&0&1\end{array}\right]$
$\mu(b)$ $=\left[\begin{array}[]{ccc}0&1&0\\ 1&1&0\\ 0&1&1\end{array}\right]$
$\mu(d)$ $=\left[\begin{array}[]{ccc}0&0&1\\ 0&1&1\\ 1&2&1\end{array}\right]$
$R$ $=\left[\begin{array}[]{c}1\\ 0\\ 0\\ \end{array}\right].$
Then $C(u)=L\mu(u)R$.
Proof.
We first obtain the following relations, for $v\in\{b,d\}^{+}$:
$C(vbb)$ $=C(v)+C(vb)+C(vd)$
$C(vbd)$ $=C(vd)$
$C(vdb)$ $=C(vb)+2C(vd)$
$C(vdd)$ $=C(v)+C(vb)+C(vd).$
We just show how to obtain the first relation; the others are similar. Letting $u=vbb$, and using (4), we get
$C(u)$ $=C(vbb)=F_{i+2j-1}+C(vb)$
$C(vb)$ $=F_{i+2j-2}+C(v)$
from which it follows that $C(u)=F_{i+2j}+C(v)$. On the other hand, from (4), we have $C(vd)=F_{i+2j-1}-C(v)$ and so
$C(v)+C(vb)+C(vd)=C(v)+F_{i+2j-2}+C(v)+F_{i+2j-1}-C(v)=C(v)+F_{i+2j},$
and the first identity follows.
Once we have the relations, we can combine them in the usual way to obtain the given linear representation. ∎
14 Going further
It is possible to study many similar recurrences using exactly the same techniques. We list just three possibilities; many more should occur to the interested reader.
  1. 1.
    One can study a variation of the defining recurrence (1) where the Lucas numbers replace the Fibonacci numbers. The resulting sequence is A118287 in the OEIS, and is computed by a $102$-state synchronized automaton.
    With this automaton one can easily prove observations such as
    A118287$(n)$ = A118287$(n-2)$ if and only if $n$ belongs to A003231.
  2. 2.
    One can generalize (1) by introducing two integer parameters, $x$ and $y$, as follows:
    $a_{x,y}(n)=\begin{cases}n,&\text{if $n\leq 1$};\\ F_{j+1}x-a_{x,y}(n-F_{j})y,&\text{if $F_{j}<n\leq F_{j+1}$ for $j\geq 2$}.\end{cases}$ (5)
    Equation (1) corresponds to $(x,y)=(1,1)$. Numerical experiments suggest that if $y=1$, the resulting sequence might be Fibonacci-synchronized for each $x\geq 1$.
    • Consider the particular case $(x,y)=(2,1)$. Then one can prove that $a_{2,1}(n)$ is computed by a $22$-state synchronized automaton. Furthermore, $((a_{2,1}(n)-1)/2)_{n\geq 1}$ is sequence A343647 and is a permutation of the positive integers.
    • The case $(x,y)=(3,2)$ seems harder to analyze. It is quite possible that it is not Fibonacci-synchronized.
  3. 3.
    Another variation is the following:
    $b(n)=\begin{cases}n,&\text{if $n\leq 1$};\\ F_{j+1}-b(b(n-F_{j})),&\text{if $F_{j}<n\leq F_{j+1}$ for $j\geq 2$}.\end{cases}$ (6)
    This sequence is generated by a $24$-state synchronized automaton.